Jebediah:

Two equations are given below: a – 3b = 9 a = b – 3 What is the solution to the set of equations in the form (a, b)? (5 points)     (–9, –6) (–4, –3) (–6, –3) (–9, –7)

DanielleLove:

you have 2 options, the easy one would be to check by replacing the variables with the numbers

Jebediah:

ok

DanielleLove:

the second one would be to actually work the equation backwards

Jebediah:

ok

dude:

You can substitute b-3 into the first equation \((b-3) – 3b = 9\) Combine like terms \(-2b-3=9\) Add 3 on both sides \(-2b=12\) Divide both sides by -2 \(b=-6\)

DanielleLove:

a – 3b = 9 a = b – 3 this equals this b-3-3b=9

dude:

Now we can substitute back into the second equation \(a = -6 – 3\\ a=-9\) So: \((-9,-6)\)

Jebediah:

oh okay

Jebediah:

thanks guys

Jebediah:

:P